the equation x(squared)=x=0 has two solutions. whats the other?
Answer by Fermat(136)

Is it supposed to be: x² + x = 0 ??
In that case,
x² + x = 0
x(x + 1) = 0
and the two solutions are,
x = 0, x = -1
=============